OpenStudy (anonymous):

What are 3 different triangle measurements that equal 4 square feet?

OpenStudy (campbell_st):

so the formula is ' \[A = \frac{1}{2} base \times height\] or \[4 = \frac{1}{2} \times base \times height\] multiply both sides of the equation by 2 and you get \[8 = base \times height\] find pairs of factors that multiply to 8 eg 1 and 8 base = 1 and height = 8